OpenStudy (anonymous):

x^y=4096 what is the largest possible integer value for x-y

OpenStudy (anonymous):

you want to maximize x and minimize y here to get x-y to be largest. The values of x and y you would thus be looking for are x=4096 and y=1. x-y=4096-1=4095

OpenStudy (anonymous):

you want to maximize the difference, D=x-y. substituting the constraint x^y=4096 ==> ylnx=ln4096 ==> y=(ln4096)/lnx, D = x - ln4096/lnx. differentiate D to find the x coordinate of the maximum value for D

OpenStudy (anonymous):

dpalnc this is an sat question so it wouldnt have calculus but thx anyway

OpenStudy (anonymous):

oopss sorry... :)

OpenStudy (anonymous):

np :P